Applied and Computational Harmonic Analysis

ACADEMIC PRESS, INC. A Division of Harcourt Brace & Company 525 B Street, Suite 1900, San Diego, California 92101-4495

November 7, 2002

1 Screen Display for Distribution Site

The Software Programs available through this Distribution Site are intended solely for the use of subscribers to Applied Computational and Harmonic Analysis. By accessing these Software Programs, you agree to be bound by the terms and conditions stated herein and in the particular Software Program being copied.

Unless otherwise specified in the applicable Software Program, you may only make a single copy of the Software Program and may not copy or otherwise retransmit the Software Program other than to make a single backup copy for archival purposes.

Unless otherwise specified in the applicable Software Program, you agree to use the software Program only for your research purposes and you agree to be bound by any other restrictions set forth in the applicable Software Program.

THE SOFTWARE PROGRAMS ARE LICENSED “AS IS,” WITHOUT WARRANTIES OF ANY KIND, AND ACADEMIC PRESS, ACHA, AND THE AUTHORS OF THE SOFTWARE PROGRAM DIS-CLAIM ANY AND ALL WARRANTIES, WHETHER EXPRESSED OR IMPLIED, INCLUDING, WITH-OUT LIMITATION, ANY IMPLIED WARRANTIES OF MERCHANTABILITY OR FITNESS FOR A PARTICULAR PURPOSE. USE OF THESE PROGRAMS IS AT YOUR OWN RISK.

IN NO EVENT SHALL ACADEMIC PRESS, ACHA, THE AUTHORS OF THE SOFTWARE PRO-GRAM, OR THE DISTRIBUTION SITE BE RESPONSIBLE FOR ANY LOSS OR DAMAGES ARIS-ING FROM ANY USE OF THE SOFTWARE PROGRAMS, INCLUDING ANY ACTUAL, INDIRECT, INCIDENTAL, CONSEQUENTIAL, SPECIAL, OR EXEMPLARY DAMAGES (INCLUDING, BUT NOT LIMITED TO, LOSS OF DATA, BUSINESS INTERRUPTION, OR LOST PROFITS) EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

2 Software Publication Policy

Applied and Computational Harmonic Analysis (ACHA) solicits submissions of computer software programs in machine-readable form. These will be reviewed by the Software Editor and, if accepted for publication, will be stored as a computer data file on a Software Distribution Site and made available by anonymous file transfer protocol (FTP)

The Software Editor is Mladen Victor Wickerhauser, VICTOR@MATH.WUSTL.EDU, Department of Mathematics, Washington University, Saint Louis, MO 63130 U.S.A. The Software Distribution Site is the home folder of user acha on WWW.MATH.WUSTL.EDU, a host computer at Washington University in Saint Louis.

2.1 Source or executable

Two types of software submissions will be considered:

• short machine-readable text files of source programs embedded in a regular article or letter for ACHA, or longer text files containing source programs which are described in a regular article or letter but are too lengthy to print in the paper journal.

• executable files, described in a regular article or letter accepted for publication in ACHA. Such submissions are appropriate for software intended to be used with computer systems lacking compilers or other development tools.

2.2 Means of publication

Software accepted for publication will be placed in a subfolder of the Software Distribution Site, labeled by the volume and number of the issue of ACHA in which an accompanying article appears.

2.3 Criteria for acceptance

Software must meet three criteria:

• Its purpose must be clear.

• It should be well documented.

• It must be properly debugged.

Authors should submit software that furthers the research and educational goals of ACHA. Software not meeting these criteria may be returned for revisions or rejected, even if the accompanying article is accepted for publication.

2.4 Transmittal

Machine-readable software should be submitted separately to the Software Editor at the same time that the accompanying article or letter is submitted to Academic Press. Once the article or letter plus the software has been reviewed and accepted for publication, the machine-readable files will be posted to the software archive so as to appear at approximately the same time as the corresponding issue of the journal.

Three methods of submitting software may be used:

• Source files containing text only may be sent by electronic mail to the Software Editor. The mail must contain the name of the author or authors and the title of the accompanying article or letter.

• Files of any sort may be uploaded to the FTP site and placed in the submissions subfolder of the Software Distribution Site. Such a submission must be accompanied by email to the Software Editor.

• Software may be submitted on floppy disk, magnetic tape, or other common physical media and sent to the Software Editor. The medium must be labeled with the author’s name, the title of the accompanying paper, the format in which the data are written such as DOS, UNIX tar(1), or Macintosh, and the date. Media will not be returned.

The author or authors should supply any information needed to unpack and use the submitted software, as well as an email address and telephone number. Questions about which data formats are acceptable should be directed to the Software Editor.

2.5 Copyrights, patents, and export restrictions

The author or authors retain the copyright to any software accepted for FTP publication, but must grant Academic Press written permission to distribute it by anonymous file transfer in accordance with the terms of the release form described below. Acceptance and posting of software on the FTP site constitute publication for patent disclosure purposes. Neither the Software Editor nor the institution maintaining the FTP site nor Academic Press will assume any responsibility for complying with applicable export laws. Authors should assume that posted software is exported to all sites with InterNet access.

2.6 Warranties and liability

Neither the Software Editor, the institution at which the FTP site is operated, ACHA, nor Academic Press will assume any liability for, and the author will indemnify these parties against any claim, loss, damage, or liability arising from, any loss, misappropriation, misuse, unauthorized use, miscommunication, or other erroneous transmission of the software.

By submitting the software to ACHA, the author is warranting that he or she has the right to authorize ACHA and its agents to use, copy, publish, and distribute the software and to sublicense such rights and that the software does not contain any materials whose use or publication violates any copyright, patent, trade secret, or other personal or proprietary right. In addition, the author warrants that the software will perform in the intended manner and will be free from any viruses or other similar defects.

2.7 Release policy

Before the software is placed on the FTP site, the authors must sign and return the ACHA Software Publication Release Form. This is available from the ACHA Editorial Office or the Software Editor or it may be downloaded from the Software Distribution Site.

3 Software Publication Release Form

Software Managing Editor: Mladen Victor Wickerhauser, VICTOR@MATH.WUSTL.EDU, Department of Mathematics, Washington University, Saint Louis, MO 63130 U.S.A.

Software Distribution Site: Home folder of user acha on WWW.MATH.WUSTL.EDU, a host computer at Washington University in Saint Louis.

Before ACHA can place your Software Program in a publicly accessible file for transmission and copying, you must read, sign, and return the accompanying release form. This form establishes the terms on which ACHA will make your Software Program available and the respective rights and responsibilities of you, ACHA, our publisher, Academic Press, and the Distribution Site. By signing this Release you will be agreeing, among other things, to the following:

• ACHA will have the right to distribute the Software Program.

• You own or are otherwise authorized to license all rights in the Software Program, including all copyrights, patents, trademarks, and any other intellectual property rights.

• You will bear all responsibility for any loss, misappropriation, misuse, or erroneous transmission of the Software or failure to comply with applicable export laws.

• You guarantee that you have the right to authorize ACHA to distribute the Software Program, and that the Software is free of viruses or other harmful code. You assume all responsibility in the event of any breach of these assurances.

We are therefore asking you to sign the accompanying form and return it to ACHA, Academic Press, Journal Division, 525 B Street, Suite 1900, San Diego, CA 92101-4495, with a copy sent to the Software Editor. No Software Programs will be processed until both copies of this form have been properly returned.

RELEASE FORM

Applied and Computational Harmonic Analysis (ACHA)

Software Program Title:

Software Program Authors:

Accompanying Manuscript Title:

Accompanying Manuscript Number:

Accompanying Manuscript Authors:

You hereby confirm the assignment to Academic Press (AP) of the following rights, upon acceptance for publication of the Software Program and its Accompanying Manuscript:

• AP or its agent may place the Software Program into a publicly accessible computer file where it may be copied through anonymous network file transfer protocol (FTP).

• AP or its agent may distribute the Software Program by electronic mail to a specific list of individuals or by Usenet network news transfer to all readers of one or more news groups.

You hereby grant to AP an irrevocable, royalty-free, nonexclusive license to use, copy, publish, and distribute the Software and to sublicense such rights in connection with making the Software available as described above. You are responsible for including an appropriate copyright notice in the Software and for including any applicable restrictions on the use, copying, retransmission, modification, or adaptation of the Software by those persons accessing the Software.

You represent and warrant that you have the right to grant the rights granted to AP and its agents herein and that the Software does not contain any materials whose use or publication violates any copyright, patent, trade secret, or other personal or proprietary right. In addition, you represent and warrant that the Software will perform in the intended manner and will be free from any viruses or other similar defects. You agree to indemnify and hold AP, ACHA, the Distribution Site, and the Software Editor harmless with respect to any claim, loss, damage, or liability arising from any breach of these representations and warranties. You are aware that the posting of Software Programs may constitute publication for patent disclosure and other legal purposes.

You acknowledge that neither AP, ACHA, the institution at which the FTP site is operated (the “Distribution Site”), nor the Software Editor will assume any liability to you or any third party for, and you agree to indemnify and hold AP, ACHA, the Distribution Site, and the Software Editor harmless with respect to any claim, loss, damage, or liability arising from, any loss, misappropriation, misuse, unauthorized use, miscommunication, or other erroneous transmission of the Software or any failure to comply with applicable export laws. You should assume that your Software Program will be exported to all sites with InterNet access.

You further acknowledge that AP and the Software Editor of ACHA are relying on this document in publishing the Software Program and that you have read and understood the Software Policy of ACHA, which is incorporated herein by reference.

Signed: Date:

Print name and title:

This form is to be signed by all the authors of the Software Program, or by an author who has obtained the written assent of the nonsigning authors, if any. In the case of a “work made for hire” (a work prepared by an employee within the scope of his or her employment or commissioned as a work made for hire under a written agreement), an authorized representative of the employer or commissioner should sign.

Please sign and return this form to ACHA at Academic Press, Journal Division, 525 B Street, Suite 1900, San Diego, CA 92101-4495, and send a copy of the signed form to the Software Editor of ACHA at the address listed above.